if line a equal 1", line b equals 1 1/2", and line c equals 2"
what does x equal if a/x = b/c
and if a/x = x/b

Answer by stanbon(75887)

If line a equal 1", line b equals 1 1/2", and line c equals 2"
what does x equal if a/x = b/c
1/x= (3/2)/2
1/x=3/4
Invert both sides to get:
x=4/3
------------



and if a/x = x/b
x^2=ab
x^2=1(3/2)
x^2=3/2
x=+(sqrt6)/2 or -(sqrt6)/2
Cheers,
Stan H.